OpenStudy (anonymous):

how do I graph a line when I have the slope and only x coordinates? Ex. Slope of 8, with points (0.5,y), (1,y),(2,y),(3,y) is that even possible or am I reading the problem wrong??

OpenStudy (amistre64):

(1,y ) (2,y) -1-y -1-y ----------- 0,0 1,0; you cant, unless you means each y is different

OpenStudy (amistre64):

then it would be (1,f(1)) (2,f(2)) -1-f(1) -1-f(1) -------------- 0 1,f(1)-f(2) = slope, y/x f(1)-f(2) = 8

OpenStudy (amistre64):

without knowing the actual function that relates x to y, there isnt a way that i can see

OpenStudy (amistre64):

(1,0) (2,8) is just as valid as (1,23) (2,31)

OpenStudy (anonymous):

all the y points are supposed to be different :S

OpenStudy (amistre64):



OpenStudy (amistre64):

you need at least one instance of an initial condition

OpenStudy (anonymous):

thats what I thought but that's all the problem gives me

OpenStudy (amistre64):

(0.5,f(1)-4) (1,f(1)) (2,f(1)+8) (3,f(1)+16)

OpenStudy (anonymous):

you that the slope\[= \tan \theta\]where \[\theta\] = the positive angle between the line and the x axis. here\[\tan \theta =8 \] => \[\theta =\tan^{-1} 8\] so draw a line making an angle so........... and then adjust it in such a way that he x coordinates satisfy the line

OpenStudy (amistre64):

dhash, there are infinte solutions without knowing how x affects y

OpenStudy (amistre64):

how can we locate a y intercept when there is no relation to associate with it?

OpenStudy (anonymous):

i guess it's just my textbook being stupid again ill have to talk to my teacher

OpenStudy (anonymous):

thanks both of you for trying tho

OpenStudy (amistre64):

good luck :)